OpenStudy (anonymous):

Solving an inequality. x^2 >= x. How to choose right domain set of solutions, please, from the equations provided.

jhonyy9 (jhonyy9):

divide both sides by x then will get x>= 1

OpenStudy (anonymous):

x will be greater than two values in this case.

OpenStudy (anonymous):

so you have \[x^2 \ge x\] \[x^2 - x \ge 0\] \[x ( x -1 ) \ge 0\] From this the book concludes that: \[x \ge 1\] or \[0 \ge x\]

OpenStudy (anonymous):

I am not sure how it derives the solution set. Does it follow like this? \[(x - 1) \ge 0\] \[x \ge 0\] ? I am not sure how they get the solution \[0 \ge x\]

jhonyy9 (jhonyy9):

yes for x<=0 this is right to

jhonyy9 (jhonyy9):

this is right too

OpenStudy (anonymous):

How do you prove that mathematically from the equations?

jhonyy9 (jhonyy9):

-1^2 >= -1 true yes ?

jhonyy9 (jhonyy9):

for zero are equale both sides

jhonyy9 (jhonyy9):

for -2^2 >= -2 4 > -2

jhonyy9 (jhonyy9):

ok ?